OpenStudy (anonymous):

7/9 + 1/6 divided by 3/7 is it 7/6?

OpenStudy (anonymous):

Is it (7/9 + 1/6)/ (3/7) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, if the expression is stated using parenthesis as shown below: \[(7/9) + ((1/6) / (3/7)) \]

OpenStudy (anonymous):

no parenthesis sorry just numbers

OpenStudy (anonymous):

It does not become 7/6 if you add 7/9 and 1/6 first.

OpenStudy (anonymous):

You must do parentheses first. So. 7/9 + 1/6 = 14/18 + 3/18 = 17/18. Then you take (17/18) / (3/7) when you divide by a fraction, the rule is to multiply by the reciprocal (flipping it upside down). So the problem becomes (17/18) * (7/3) so you get 119/54 then you may want it as a mixed number.. 2 11/54